# GMoE: Empowering LLMs Fine-Tuning via MoE Graph Collaboration

### v3

@@ Abstract
The sparse Mixture-of-Experts (MoE) architecture of large language models (LLMs) confronts an inherent issue of load imbalance arising from the simplistic linear router strategy, which ultimately causes the instability and inefficient learning of LLMs. To address this challenge, we introduce a novel MoE graph-based framework GMoE, aimed at enhancing the collaboration among multiple experts. In GMoE, a graph router function is designed to capture the collaboration signals among experts. This enables all experts to dynamically allocate information derived from input data by sharing information with their neighboring experts. Moreover, we put forward two coordination strategies in GMoE: the Poisson distribution-based distinction strategy and the Normal distribution-based balance strategy, to further release the capacity of each expert and increase the model stability in the fine-tuning of LLMs. Specifically, we leverage a parameter-efficient fine-tuning technique, i.e., Low-Rank Adaptation (LoRA), to implement the graph MoE architecture. Extensive experiments on four real-world benchmark datasets demonstrate the effectiveness of GMoE, showing the benefits of facilitating collaborations of multiple experts in LLM fine-tuning. The code of experimental implementation is available at https://github.com/BAI-LAB/GMoE
@@
Zenan Xu Tencent AI Lab zenanxu@tencent.com                        Zhe Zhao Tencent AI Lab nlpzhaozhe@tencent.com                        Chuan Shi Beijing University of Posts and Telecommunication shichuan@bupt.edu.cn
@@ 1 Introduction
The Mixture-of-Expert (MoE) architecture has emerged as a promising approach to enhance the overall performance of large language models (LLMs) under the scaling law theory. The application of MoE in the fine-tuning of LLMs has drawn significant attention due to the substantial improvements it brings to model performance in practical downstream applications. In a typical MoE setup, a simple linear function, acting as the “router function” or “gating function,” assigns weights to each expert based on the information from input tokens, then the top-k experts with largest assigned weights are activated in the fine-tuning process. However, sparse MoE methods face significant challenges that impede their performance and introduce model instability, with a core issue being their linear router strategy that creates severe load imbalance, where a few experts are over-trained while others are under-utilized Fedus et al. (2022).
Recent studies have introduced load balance losses to constrain expert allocation frequencies, thereby alleviating MoE load imbalance (Fedus et al., 2022; Shazeer et al., 2016; Zoph et al., 2022; Luo et al., 2024; Dou et al., 2023a; Wang et al., 2022). However, these approaches primarily focus on regularization-based constraints for load balancing while failing to address a critical limitation: the absence of effective communication and collaborative mechanisms among experts during router allocation. Without expert collaboration, the router relies solely on individual input patterns instead of leveraging collective capabilities. This coordination gap exacerbates inefficiencies: even with balanced allocations, the lack of inter-expert communication prevents dynamic optimization to exploit each expert’s unique strengths. As a result, experts remain underutilized, limiting the model’s potential and leading to instability during LLM fine-tuning.
@@ 2.1 Mixture-of-Expert (MoE)
Empowered by the collaboration of multiple experts, the Mixture-of-Expert (MoE) framework has achieved remarkable performance in many applications (Chen et al., 2024; Li et al., 2024a; Lin et al., 2024; Li et al., 2024c; Zadouri et al., 2023). Each expert in the MoE framework specializes in handling a subset of input information and different experts coordinate together to gain benefits for the overall MoE framework. According to the number of experts activated by the router function, the MoE methods can be classified into two categories, i.e., dense MoE and sparse MoE (Dou et al., 2023b; Li et al., 2024b). In dense MoE, all experts are activated in the learning process, which enhances the model capability but suffers from high computational costs. In sparse MoE, only a selected subset of experts are activated to use their specialized knowledge and achieve optimal results, thereby reducing computational overhead. The participation of each expert in the computation process is assigned by a router function (or gating function), ensuring an optimal blend of their specialized contributions (Zoph et al., 2022; Tang et al., 2025). Specifically, the router is usually a simple linear function that controls the engagement of expert computations. In the dense MoE, all experts are allocated to computation depending on the weights assigned by the router. In sparse MoE, only Top-K experts with the highest weights are selectively activated in the learning process (Shazeer et al., 2016).
For example, the typical sparse MoE approach Switch Transformers (Fedus et al., 2022) routes only a single expert for the input token to reduce the computation costs, but it faces the load imbalance problem due to lacking collaboration of multiple experts. Existing studies have incorporated auxiliary losses (Bengio et al., 2015; Fedus et al., 2022) to maintain a balanced load among experts in sparse MoE architectures. However, exploring an optimal collaboration mechanism among the activated experts to enhance their load distribution remains an ongoing challenge in this field. In our work, we propose a novel graph router to assign weights to experts with consideration of their collaborative interactions on the MoE Graph, so as to achieve better capability and stability in fine-tuning of LLMs.
@@ 2.2 PEFT with MoE
The aim of fine-tuning LLMs is to optimize the model’s performance in specific downstream tasks. Due to the high computational costs of updating all parameters in the full fine-tuning approach, the Parameter-Efficient Fine-Tuning (PEFT) technique is proposed to address this problem (Hu et al., 2021; Houlsby et al., 2019; Li and Liang, 2021; Lester et al., 2021; Tian et al., 2024). In PEFT, only a small subset of parameters are updated of the base LLMs, making it more efficient and practical in real applications. The most widely used PEFT technique is LoRA (Hu et al., 2021), which uses a low-rank decomposition technique and updates the decomposed parameter matrix in the model training.
Recent studies have shown improvements of model performance by integrating the MoE framework with PEFT in LLMs fine-tuning literature (Wu et al., 2024; Zadouri et al., 2023). In the transformer block of LLMs, each expert is a LoRA module working on either the FFN layer (Dou et al., 2023b; Wang et al., 2022; Li et al., 2024a), attention layer (Luo et al., 2024; Gou et al., 2023; Zhu et al., 2023), the whole transformer block (Zadouri et al., 2023; Liu et al., 2023; Gao et al., 2024) and each layer (Wu et al., 2024). Then a router function of MoE is designed to blend the contributions of all experts. Among them, the adoption of LoRA in the FFN layer attracted the most attention due to its extensive applicability in various LLM tasks in recent years.
Our work focuses on exploring the effective MoE collaboration mechanism to enhance the PEFT of LLMs in downstream tasks. The MoE consists of multiple LoRA components that work on the FFN layer in the transformer block. Different from the conventional router function employed in existing MoE studies, where the weight assigned to each expert is solely dependent on its own input information, we design a novel graph router, which leverages graph neural networks to learn the collaboration information among all experts based on a MoE graph.
@@ 3.1 Low-Rank Adaptation (LoRA)
Low-rank adaptation (LoRA) (Hu et al., 2021) is a widely used parameter-efficient fine-tuning (PEFT) technique in pre-trained LLMs. LoRA adopts a low-rank decomposition technique to update the parameters of the decomposed parameter matrix to learn the data distribution of the specific downstream task. It works on the feed-forward layer (FFN layer) in a transformer block of LLMs. For a linear layer in FFN represented as $\mathbf{h}=\mathbf{W}\mathbf{x}$, the update of the decomposed parameter matrix of LoRA is defined as:
@@ 3.2 Mixture-of-Expert (MoE-LoRA)
where $\mathcal{E}_{i}$ is the $i$th expert network and $N$ is the number of experts. The router function $\mathcal{R}(\cdot)$ employs a linear function to produce a probability distribution from input information, which serves as the weights allocated to all experts. Following the sparse-gated strategy in MoE studies (Zoph et al., 2022; Li et al., 2024a), which had been proposed to address the computational overhead problem in LLMs. The top-K experts with the highest weights are selectively activated.
@@ 4.1 Graph Router
The router in GMoE takes full advantage of the collaborative information aggregations by GNNs from other expert nodes. This alleviates the load unbalance and localized convergence tendencies caused by a small number of experts activated in the existing MoE studies (Shazeer et al., 2016). Except for introducing a collaboration mechanism of all experts, we use two coordination strategies, i.e., the Poisson distribution-based expert distinction strategy and the normal distribution-based load balance strategy in GMoE, to further release the capacity of each expert and increase the model stability.
@@ 4.2 Expert Distinction Strategy
[…]
@@ 4.3 Load Balance Strategy
Apart from the distinct capabilities of different experts, one important factor in MoE training is keeping the load balance of all experts. Otherwise, greater routing weights on a small number of experts in the early stages of the fine-tuning process will result in a rapid localized optimization problem. In GMoE, we propose the Normal distribution-based load balance strategy. Specifically, as only Top-K experts are activated in dealing with input information, we calculate the cumulative weights of each expert to represent its activation frequency. Then construct the activation frequency vector of all experts by normalizing the cumulative weights, denoted as $\mathbf{v}_{a}$ in the next feed-forward step. Our aim is to make the activation frequency vector $\mathbf{v}_{a}$ follow a natural normal distribution rather than the absolute equality in existing MoE literature (Li et al., 2024a; Zoph et al., 2022). Formally, the Normal distribution-based load balance loss function can be defined as:
@@
[…]
@@ 5.1 Experimental Settings
To evaluate the performance of our framework, we use four representative public datasets, including the question-answering task, i.e., ARC-Challenge (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), SIQA (Sap et al., 2019), and task classification task in BoolQ (Clark et al., 2019) dataset. These datasets cover the evaluations on different domains of LLMs, such as factual knowledge from Wikipedia, natural science, science facts, and social interactions.
@@
To verify the effectiveness of GMoE, we compare the performance of typical PEFT MoE methods on three popular open-source LLMs, i.e., Llama3-8B [^1] https://github.com/meta-llama/llama3, Qwen2-7B (Yang et al., 2024), and Yi-1.5-9B (Young et al., 2024).
LoRAMoE (Dou et al., 2023b).
@@
All methods are implemented in PyTorch with an A800-80G. For each baseline method, a grid search is applied to find the optimal settings. These include learning rate from $\{0.1,0.01,0.001,0.0001,0.00001\}$, number of experts from $\{4,8,12,16\}$, rank of LoRA from $\{1,2,4,8,16\}$, Top-K experts from $\{1,2,3,4,5\}$. We report the results of each method with its optimal hyperparameter settings on the validation data. In our model, we adopt GCN as the graph neural network (GNN). The number of aggregation layers in GCN is 2, and the hidden dimension of GCN layers is 256. The edge density $\beta$ to construct the MoE graph is $0.1$. The $\alpha$ in Eq. 1 is 4 in the LoRA component. The coefficients of the Poisson distribution loss function and Normal distribution loss function are set to 0.005 and 8 in the final loss function in LLM fine-tuning.

### v4

@@ Abstract
The sparse Mixture-of-Experts (MoE) architecture of large language models (LLMs) confronts an inherent issue of load imbalance arising from the simplistic linear router strategy, which ultimately causes the instability and inefficient learning of LLMs. To address this challenge, we introduce a novel MoE graph-based framework GMoE, aimed at enhancing the collaboration among multiple experts. In GMoE, a graph router function is designed to capture the collaboration signals among experts. This enables all experts to dynamically allocate information derived from input data by sharing information with their neighboring experts. Moreover, we put forward two coordination strategies in GMoE: the Poisson distribution-based distinction strategy and the Normal distribution-based balance strategy, to further release the capacity of each expert and increase the model stability in the fine-tuning of LLMs. Specifically, we leverage a parameter-efficient fine-tuning technique, i.e., Low-Rank Adaptation (LoRA), to implement the graph MoE architecture. Extensive experiments on four real-world benchmark datasets demonstrate the effectiveness of GMoE, showing the benefits of facilitating collaborations of multiple experts in LLM fine-tuning. The code of experimental implementation is available at https://github.com/BAI-LAB/GMoE.
@@
Zenan Xu xuzn@alumni.sysu.edu.cn          Chuan Shi Beijing University of Posts and Telecommunication shichuan@bupt.edu.cn
@@ 1 Introduction
The Mixture-of-Expert (MoE) architecture has emerged as a promising approach to enhance the overall performance of large language models (LLMs) under the scaling law theory. The application of MoE in the fine-tuning of LLMs has drawn significant attention due to the substantial improvements it brings to model performance in practical downstream applications. In a typical MoE setup, a simple linear function, acting as the “router function” or “gating function,” assigns weights to each expert based on the information from input tokens, then the top-k experts with largest assigned weights are activated in the fine-tuning process. However, sparse MoE methods face significant challenges that impede their performance and introduce model instability, with a core issue being their linear router strategy that creates severe load imbalance, where a few experts are over-trained while others are under-utilized switch-trans.
Recent studies have introduced load balance losses to constrain expert allocation frequencies, thereby alleviating MoE load imbalance (switch-trans; loadbal; moe20222; moeloracontra; loramoe4; wang2022adamix). However, these approaches primarily focus on regularization-based constraints for load balancing while failing to address a critical limitation: the absence of effective communication and collaborative mechanisms among experts during router allocation. Without expert collaboration, the router relies solely on individual input patterns instead of leveraging collective capabilities. This coordination gap exacerbates inefficiencies: even with balanced allocations, the lack of inter-expert communication prevents dynamic optimization to exploit each expert’s unique strengths. As a result, experts remain underutilized, limiting the model’s potential and leading to instability during LLM fine-tuning.
@@ 2.1 Mixture-of-Expert (MoE)
Empowered by the collaboration of multiple experts, the Mixture-of-Expert (MoE) framework has achieved remarkable performance in many applications (chen2024llava; li2024mixlora; lin2024moe; li2024uni; zadouri2023pushing). Each expert in the MoE framework specializes in handling a subset of input information and different experts coordinate together to gain benefits for the overall MoE framework. According to the number of experts activated by the router function, the MoE methods can be classified into two categories, i.e., dense MoE and sparse MoE (loramoe; li2024locmoe). In dense MoE, all experts are activated in the learning process, which enhances the model capability but suffers from high computational costs. In sparse MoE, only a selected subset of experts are activated to use their specialized knowledge and achieve optimal results, thereby reducing computational overhead. The participation of each expert in the computation process is assigned by a router function (or gating function), ensuring an optimal blend of their specialized contributions (moe20222; tang2025graphmoe). Specifically, the router is usually a simple linear function that controls the engagement of expert computations. In the dense MoE, all experts are allocated to computation depending on the weights assigned by the router. In sparse MoE, only Top-K experts with the highest weights are selectively activated in the learning process (loadbal).
For example, the typical sparse MoE approach Switch Transformers (switch-trans) routes only a single expert for the input token to reduce the computation costs, but it faces the load imbalance problem due to lacking collaboration of multiple experts. Existing studies have incorporated auxiliary losses (bengio2015conditional; switch-trans) to maintain a balanced load among experts in sparse MoE architectures. However, exploring an optimal collaboration mechanism among the activated experts to enhance their load distribution remains an ongoing challenge in this field. In our work, we propose a novel graph router to assign weights to experts with consideration of their collaborative interactions on the MoE Graph, so as to achieve better capability and stability in fine-tuning of LLMs.
@@ 2.2 PEFT with MoE
The aim of fine-tuning LLMs is to optimize the model’s performance in specific downstream tasks. Due to the high computational costs of updating all parameters in the full fine-tuning approach, the Parameter-Efficient Fine-Tuning (PEFT) technique is proposed to address this problem (lora; houlsby2019parameter; li2021prefix; lester2021power; tian2024hydralora). In PEFT, only a small subset of parameters are updated of the base LLMs, making it more efficient and practical in real applications. The most widely used PEFT technique is LoRA (lora), which uses a low-rank decomposition technique and updates the decomposed parameter matrix in the model training.
Recent studies have shown improvements of model performance by integrating the MoE framework with PEFT in LLMs fine-tuning literature (wu2024mixture; zadouri2023pushing). In the transformer block of LLMs, each expert is a LoRA module working on either the FFN layer (loramoe; wang2022adamix; li2024mixlora), attention layer (moeloracontra; loramoe2; zhu2023sira), the whole transformer block (zadouri2023pushing; loramoe3; mola) and each layer (wu2024mixture). Then a router function of MoE is designed to blend the contributions of all experts. Among them, the adoption of LoRA in the FFN layer attracted the most attention due to its extensive applicability in various LLM tasks in recent years.
Our work focuses on exploring the effective MoE collaboration mechanism to enhance the PEFT of LLMs in downstream tasks. The MoE consists of multiple LoRA components that work on the FFN layer in the transformer block. Different from the conventional router function employed in existing MoE studies, where the weight assigned to each expert is solely dependent on its own input information, z we design a novel graph router, which leverages graph neural networks to learn the collaboration information among all experts based on a MoE graph.
@@ 3.1 Low-Rank Adaptation (LoRA)
Low-rank adaptation (LoRA) (lora) is a widely used parameter-efficient fine-tuning (PEFT) technique in pre-trained LLMs. LoRA adopts a low-rank decomposition technique to update the parameters of the decomposed parameter matrix to learn the data distribution of the specific downstream task. It works on the feed-forward layer (FFN layer) in a transformer block of LLMs. For a linear layer in FFN represented as $\mathbf{h}=\mathbf{W}\mathbf{x}$, the update of the decomposed parameter matrix of LoRA is defined as:
@@ 3.2 Mixture-of-Expert (MoE-LoRA)
where $\mathcal{E}_{i}$ is the $i$th expert network and $N$ is the number of experts. The router function $\mathcal{R}(\cdot)$ employs a linear function to produce a probability distribution from input information, which serves as the weights allocated to all experts. Following the sparse-gated strategy in MoE studies (moe20222; li2024mixlora), which had been proposed to address the computational overhead problem in LLMs. The top-K experts with the highest weights are selectively activated.
@@ 4.1 Graph Router
The router in GMoE takes full advantage of the collaborative information aggregations by GNNs from other expert nodes. This alleviates the load unbalance and localized convergence tendencies caused by a small number of experts activated in the existing MoE studies (loadbal). Except for introducing a collaboration mechanism of all experts, we use two coordination strategies, i.e., the Poisson distribution-based expert distinction strategy and the normal distribution-based load balance strategy in GMoE, to further release the capacity of each expert and increase the model stability.
@@ 4.2 Expert Distinction Strategy
$\textit{Loss-Poisson}=\sum_{i}^{N}\mathbf{v}_{\textit{poisson}(\lambda,i)}\log\frac{\mathbf{v}_{\textit{poisson}(\lambda,i)}}{{\mathbf{v}_{r}}},$ (7)
@@ 4.3 Load Balance Strategy
Apart from the distinct capabilities of different experts, one important factor in MoE training is keeping the load balance of all experts. Otherwise, greater routing weights on a small number of experts in the early stages of the fine-tuning process will result in a rapid localized optimization problem. In GMoE, we propose the Normal distribution-based load balance strategy. Specifically, as only Top-K experts are activated in dealing with input information, we calculate the cumulative weights of each expert to represent its activation frequency. Then construct the activation frequency vector of all experts by normalizing the cumulative weights, denoted as $\mathbf{v}_{a}$ in the next feed-forward step. Our aim is to make the activation frequency vector $\mathbf{v}_{a}$ follow a natural normal distribution rather than the absolute equality in existing MoE literature (li2024mixlora; moe20222). Formally, the Normal distribution-based load balance loss function can be defined as:
@@
$\textit{Loss-Normal}=\sum_{i}^{N}\mathbf{v}_{\textit{normal}(\mu,\sigma,i)}\log\frac{\mathbf{v}_{\textit{normal}(\mu,\sigma,i)}}{{\mathbf{v}_{a}}},$ (9)
@@ 5.1 Experimental Settings
To evaluate the performance of our framework, we use four representative public datasets, including the question-answering task, i.e., ARC-Challenge (arc), OpenBookQA (obqa), SIQA (siqa), and task classification task in BoolQ (boolq) dataset. These datasets cover the evaluations on different domains of LLMs, such as factual knowledge from Wikipedia, natural science, science facts, and social interactions.
@@
To verify the effectiveness of GMoE, we compare the performance of typical PEFT MoE methods on three popular open-source LLMs, i.e., Llama3-8B [^1] https://github.com/meta-llama/llama3, Qwen2-7B (qwen2), and Yi-1.5-9B (yi).
LoRAMoE (loramoe).
@@
For each baseline method, a grid search is applied to find the optimal settings. These include learning rate from $\{0.1,0.01,0.001,0.0001,0.00001\}$, number of experts from $\{4,8,12,16\}$, rank of LoRA from $\{1,2,4,8,16\}$, Top-K experts from $\{1,2,3,4,5\}$. We report the results of each method with its optimal hyperparameter settings on the validation data. In our model, we adopt GCN as the graph neural network (GNN). The number of aggregation layers in GCN is 2, and the hidden dimension of GCN layers is 256. The edge density $\beta$ to construct the MoE graph is $0.1$. The $\alpha$ in Eq. 1 is 4 in the LoRA component. The coefficients of the Poisson distribution loss function and Normal distribution loss function are set to 0.005 and 8 in the final loss function in LLM fine-tuning.